# Age-Based Scheduling for Mobile Edge Computing: A Deep Reinforcement Learning Approach

### 01 Dec 2023 v1

@@ VI-B In Comparison with Benchmark Algorithms
To demonstrate the benefits of PDS, we also implement the average-reward DDPG (A-DDPG) presented in Section V-A and the conventional discounted-reward DDPG (D-DDPG). Since the convergence speed of both algorithms is slow, we have extended the duration of the experiment to $10^{5}$ time slots. As shown in Fig. 8, the AoI for both algorithms continues to increase to very large values. Although A-DDPG’s AoI increases relatively slowly, its average energy consumption is much higher than the budget. Additionally, we have also observed that the two algorithms are particularly sensitive to neural architecture and parameter settings, making them prone to training instability. Therefore, for our problem, using DDPG directly cannot yield an effective scheduling policy within an acceptable time.
@@ VI-C Performance under Different Numbers of WDs
Fig. 9 shows the average AoI of different algorithms under different numbers of WDs. Notice that the energy budgets of COO and LPO are still two times and three times that of DPDS. In our problem, since the total spectrum bandwidth of the BS is fixed, having more WDs means that each WD can occupy fewer spectrum resources. According to Shannon’s formula, the wireless transmission rate is linear to the spectrum bandwidth. Therefore, for scheduling schemes like COO, which can only perform computation offloading, the number of WDs has a significant impact on AoI. In contrast, the average AoI of LPO remains the same regardless of the number of WDs. For DPDS and DPL, since they partially rely on computation offloading, the increase in the number of WDs also leads to an increase in AoI, but the overall impact is much smaller than that of COO.
@@ VI-D Performance under Different Energy Budgets
To analyze the impact of the energy consumption constraint on AoI, we conduct experiments under different energy budgets. Specifically, we gradually increase the energy budget of each algorithm with a multiplier. Therefore, the energy budgets for COO and LPO are consistently two times and three times that of DPDS. Our experiment results are displayed in Fig. 10. As expected, the AoI for all algorithms decreases with the increase in the energy budget, but the rate of decrease continues to diminish. Among the four algorithms, COO is the most sensitive to the energy budget, especially when the energy budget is relatively low. For example, when we initially increase the energy budget by $10\%$, the average AoI of COO decreases from $23.1$ to $2.8$, showing an improvement of almost an order of magnitude. Combining with the data presented in the previous subsection, we can slightly increase the energy budget when the number of WDs increases, thus offsetting the negative impact of insufficient spectrum bandwidth on AoI.
@@ VII Conclusion
In this paper, we have delved into age-based scheduling for real-time applications in MEC systems. Initially, we have refined the concept of AoI to accommodate the event-driven sampling policy and additional processing time. Subsequently, we have formulated the problem of AoI minimization and transformed it into an equivalent CMDP. We have demonstrated that CMDP can be effectively tackled using RL methods and introduced the novel concept of PDS to accelerate RL’s learning process. In order to address the limitations inherent in tabular RL, we have merged our algorithm with DDPG and proposed a deep PDS learning approach. Numerical results underscore the remarkable efficiency of our algorithm, showcasing its superior performance against benchmark methods in diverse scenarios. As we look to the future, we aspire to (i) extend the applicability of the deep PDS learning algorithm to a wider array of scenarios and (ii) enhance its performance by refining neural network architecture and training strategies.

### 23 Feb 2024 v2

@@ VI-B In Comparison with Benchmark Algorithms
The selection of LPO and COO as benchmark algorithms is motivated by their frequent use in related literature studying computation offloading in MEC systems [21, 22, 34, 69, 70]. In particular, they serve as common baselines to highlight the advantages due to the strategic combination of local computing and computation offloading.
@@
The proposed DPDS algorithm can be regarded as a fusion of PDS and DDPG. For comparison purpose, we implement the average-reward DDPG (A-DDPG) presented in Section V-A and the conventional discounted-reward DDPG (D-DDPG). Since the convergence speed of both algorithms is slow, we have extended the duration of the experiment to $10^{6}$ time slots. However, even after such a long period, the algorithms failed to converge to an effective scheduling policy. As shown in Fig. 8, the AoI for both algorithms continues to increase to very large values. Although A-DDPG’s AoI increases relatively slowly, its average energy consumption is much higher than the budget. Additionally, we have also observed that the two algorithms are particularly sensitive to neural architecture and parameter settings, making them prone to training instability. The above observation fully demonstrates the advantages of DPDS over DDPG, which is primarily attributed to the introduction of PDS and the techniques presented in Section V-C.
@@ VI-C Performance under Different Numbers of WDs
Fig. 9 shows the average AoI of different algorithms under different numbers of WDs. Notice that the energy budgets of COO and LPO are still two times and three times that of DPDS. The rest parameters remain unchanged as described at the beginning of this section. In our problem, since the total spectrum bandwidth of the BS is fixed, having more WDs means that each WD can occupy fewer spectrum resources. According to Shannon’s formula, the wireless transmission rate is linear to the spectrum bandwidth. Therefore, for scheduling schemes like COO, which can only perform computation offloading, the number of WDs has a significant impact on AoI. In contrast, the average AoI of LPO remains the same regardless of the number of WDs. For DPDS and DPL, since they partially rely on computation offloading, the increase in the number of WDs also leads to an increase in AoI, but the overall impact is much smaller than that of COO.
@@ VI-D Performance under Different Energy Budgets
To analyze the impact of the energy consumption constraint on AoI, we conduct experiments under different energy budgets. Specifically, we gradually increase the energy budget of each algorithm with a multiplier. Therefore, the energy budgets for COO and LPO are consistently two times and three times that of DPDS. Our experiment results are displayed in Fig. 10. As expected, the AoI for all algorithms decreases with the increase in the energy budget, but the rate of decrease continues to diminish. Notice that both DPL and DPDS aim to process all queueing tasks as soon as possible. However, due to the difference in the definition of AoI and delay, the optimal scheduling policy of tasks for these two metrics is different. This explains why DPL produces sub-optimal AoI compared to DPDS. Nevertheless, as $E_{i}^{max}$ keeps growing, the processing capabilities of WDs become large enough so that almost all tasks can be completed immediately. In this case, the scheduling policy of both algorithms is very similar and results in nearly identical AoI.
Among the four algorithms, COO is the most sensitive to the energy budget, especially when the energy budget is relatively low. For example, when we initially increase the energy budget by $10\%$, the average AoI of COO decreases from $23.1$ to $2.8$, showing an improvement of almost an order of magnitude. Combining with the data presented in the previous subsection, we can slightly increase the energy budget when the number of WDs increases, thus offsetting the negative impact of insufficient spectrum bandwidth on AoI.
@@ VII Conclusion
In this paper, we have delved into age-based scheduling for real-time applications in MEC systems. Initially, we have refined the concept of AoI to accommodate the event-driven sampling policy and additional processing time. Subsequently, we have formulated the problem of AoI minimization and transformed it into an equivalent CMDP. We have demonstrated that CMDP can be effectively tackled using RL methods and introduced the novel concept of PDS to accelerate RL’s learning process. In order to address the limitations inherent in tabular RL, we have merged our algorithm with DDPG and proposed a deep PDS learning approach. Numerical results underscore the remarkable efficiency of our algorithm, showcasing its superior performance against benchmark methods in diverse scenarios. In our future work, we aspire to (i) extend the applicability of the deep PDS learning algorithm to a wider array of scenarios and (ii) enhance its performance by refining neural network architecture and training strategies.